Exercise 5.

Find the solution of

dy/dx= e2x+y 2x+y

that has y = 0 when x = 0


1
Expert's answer
2021-07-16T13:14:59-0400

Given equation,


"\dfrac{dy}{dx}=e^{2x+y}"


"\implies \dfrac{dy}{dx}=e^{2x}.e^y"


"\implies e^{-y} dy=e^{2x} dx"


Integrate Both the sides


"\implies \int e^{-y} dy=\int e^{2x} dx"



"\implies -e^{-y}=\dfrac{e^{2x}}{2}+C"



As,y(0)=0,


"\implies -e^{-0}=\dfrac{e^0}{2}+C"


So "C=-1-\dfrac{1}{2}=\dfrac{-3}{2}"



So the solution is


"-e^{-y}=\dfrac{e^{2x}}{2}-\dfrac{3}{2}"


i.e. "e^{-y}=-\dfrac{e^{2x}}{2}+\dfrac{3}{2}"